OpenStudy (anonymous):

2x2 - 4x - 3 from x2 - 5x - 8.

OpenStudy (anonymous):

2x2 - 4x - 3 from x2 - 5x - 8 x^2 - 5x - 8 - (2x^2 -4x-3) = x^2 -5x -8 - 2x^2 +4x+3 = -x^2 -x -5

OpenStudy (anonymous):

Please type the complete question ..